# Learning Diverse Robot Striking Motions with Diffusion Models and Kinematically Constrained Gradient Guidance

### v1

@@ Abstract
Advances in robot learning have enabled robots to generate skills for a variety of tasks. Yet, robot learning is typically sample inefficient, struggles to learn from data sources exhibiting varied behaviors, and does not naturally incorporate constraints. These properties are critical for fast, agile tasks such as playing table tennis. Modern techniques for learning from demonstration improve sample efficiency and scale to diverse data, but are rarely evaluated on agile tasks. In the case of reinforcement learning, achieving good performance requires training on high-fidelity simulators. To overcome these limitations, we develop a novel diffusion modeling approach that is offline, constraint-guided, and expressive of diverse agile behaviors. The key to our approach is a kinematic constraint gradient guidance (KCGG) technique that computes gradients through both the forward kinematics of the robot arm and the diffusion model to direct the sampling process. KCGG minimizes the cost of violating constraints while simultaneously keeping the sampled trajectory in-distribution of the training data. We demonstrate the effectiveness of our approach for time-critical robotic tasks by evaluating KCGG in two challenging domains: simulated air hockey and real table tennis. In simulated air hockey, we achieved a 25.4% increase in block rate, while in table tennis, we saw a 17.3% increase in success rate compared to imitation learning baselines.
@@ I Introduction
Figure 1: Overview of KCGG: Our diffusion model was trained on four kinesthetically demonstrated strokes. At inference, for each denoising step, our KCGG method computes the cost given the constraint function and intermediate trajectory. Then, the gradient with respect to the prior trajectory through the diffusion model is computed to update the sample.
In this paper, we develop a novel imitation learning based approach that enables robots to learn from small offline datasets to perform agile tasks in dynamical environments. We ground our work in application to striking motions evaluated in established challenge domains for robotics: a virtual air hockey domain [20, 9] and a physical table tennis domain [21, 22]. We show that our diffusion-based approach can learn to condition trajectories to meet environment constraints and perform tasks difficult to specify (e.g. hitting a topspin, a high lob, etc) with just 15 kinesthetic demonstrations per stroke type, which classical motion planning cannot achieve without explicit task heuristics. The key to our approach is the novel formulation of a kinematic constraint gradient guidance (KCGG) technique that enables the robot to balance between constraints and adhering to the demonstration distribution.
@@
  • We show KCGG improves a real-world table tennis robot’s ability to return ping-pong balls. Our method improves success rate by 124.5% over diffusion baselines [23, 24] and 17.3% over state of the art imitation learning baselines [19], demonstrating efficacy in agile robotic tasks.
@@ II Related Work
LfD [15, 16, 17] aims to enable robots to acquire new skills by learning to imitate an expert. MP-based LfD has been explored extensively for agile striking motions in table tennis by learning movement primitives. For example, Mixture of Movement Primitives (MoMP) [22, 25] learns to select a dynamic movement primitive (DMP) from a library of pre-trained DMPs to generate joint trajectories for returning the ball. Probabilistic movement primitives (ProMP) [26], which model the variability in movements as trajectory distributions, were adapted to generate striking motions that maximize the likelihood of ball return [27, 28, 29, 30]. Movement primitive approaches are sample-efficient; however, generating diverse motions (e.g. slice vs. lob in table tennis) is difficult without a hierarchical approach.
Recently, RL methods that are utilizing heavy computational resources are enabling systems to explore in task space without using demonstrated expert knowledge [21, 31, 32, 33, 34]. However, most of them require large training datasets [21, 32], carefully shaped reward [21], or human behavior modeling [34] which require a simulation. In contrast, we develop a fully offline method that requires few demonstrations and captures diverse agile behaviors in a single model, simplifying the training and deployment process onto real-world robots.
@@
Diffusion models [35, 36] are a class of generative models popularized in computer vision that have recently shown promise in robotics for imitation learning [23], offline RL [37, 38], and motion planning [23, 39]. Compared with previous generative modeling approaches such as GANs [40], diffusion models generally exhibit improved sample quality and diversity with a tradeoff in sampling time. In this work, we enhance the quality of samples generated by diffusion models while maintaining the same sampling time. This approach allows us to produce higher-quality motion samples without increasing computational cost, which is crucial for learning agile motions.
Recent works using diffusion models for robotic manipulation, such as BESO [41] and MPD2 [42], have demonstrated impressive performance and high success rates in challenging domains requiring millimeter-level precision. Our work, however, explores the additional complexity of solving dynamic tasks. Furthermore, while BESO [41] is designed for goal-conditioned tasks, the goal space in our domain is difficult to define. Consequently, our method does not rely on goal conditioning and instead adapts during sampling time.
One advantage of diffusion models is the ability to guide samples towards certain objectives at sampling time. Several prior works in robotics have shown how to add constraints for motion planning [24] and state estimation [43]. The computer vision community has also investigated using diffusion models to solve inverse problems [44, 45], where the goal is to reconstruct samples from sparse measurements. We take inspiration from these techniques used in inverse problems to formulate a novel method for guiding diffusion samples towards kinematics constraints. Our method is simple and enables the constraints to be better incorporated into the resulting diffusion samples.
Figure 2: Visualization of Diffusion Sampling Process in Air-Hockey Defend: Our KCCG method moves the conditioned points in the trajectory (shown in green) towards a predicted future puck position (shown as the orange point).
@@ III Problem Formulation
The agile robot striking problem can be formulated as a motion planning task, where the objective is to determine a trajectory for the robot that adheres to a variety of constraints. We define $\tau=\{s_{0},s_{1},...s_{H}\}$ as a sequence of states up to a time horizon $H$. Each state $s=[q,\dot{q}]\in\mathbb{R}^{d}$ is composed of the robot’s joint positions $q$ and joint velocities $\dot{q}$. The goal of motion planning is to identify a trajectory between initial and final states that satisfy a set of costs or constraints $c_{j}(\tau)$.
Trajectory optimization-based approaches formulate the motion-planning problem as $\mathcal{J}(\tau)=\sum_{j}\lambda_{j}c_{j}(\tau)$, where the objective is to minimize the total cost $\mathcal{J}$ while satisfying the various constraints. Other approaches formulate the motion-planning problem through probabilistic inference. Here, the optimality of a state within the trajectory is represented as $\mathcal{O}_{t}$ and the goal is to sample from a posterior distribution where the $p(\tau|\mathcal{O}_{1:t})\sim p(\tau)p(\mathcal{O}_{1:t}|\tau)$, turning the motion planning problem into a conditional sampling problem [23]. We use this formulation, where the diffusion model learns the prior $p(\tau)$, and we use the guided sampling formulation [46] to conditionally sample the model on our task constraints.
Task constraints of interest for the agile striking problem include the robot kinematic constraints and ensuring contact between the end-effector and the moving object, where $b\in\mathbb{R}^{3}$ is the object’s position in Cartesian space. The forward kinematics function that maps the robot joint positions $q$ to end-effector position $x$ is denoted as $F\colon q\rightarrow x,q\in R^{d},x\in R^{3}$. Our constraint minimizes the distance between the end-effector and object positions at timestep $t$ within a time window $[t_{s},t_{e}]$ where the robot can feasibly strike the object, i.e. $c(\tau)=\min_{t\in[t_{s},t_{e}],q_{t}\in\tau}\|F(q_{t})-b_{t}\|_{2}^{2}$. Our method is amenable to other types of constraints (e.g. for obstacle avoidance).
@@ IV Preliminaries: Score-Based Diffusion Models
Diffusion models are a class of generative models that learn to generate samples through a forward noising process and a reverse de-noising process. The forward noising process is commonly defined by: $d\tau_{t}=f(\tau_{t},t)\,dt+\sigma(t)\,dw_{t}$, where $\tau$ is the state trajectory of the robot, $f(\tau_{t},t)$ is the drift term and $\sigma(t)$ is the standard Brownian noise process. The forwards process is coupled with a backwards process: $d\tau_{t}=\left[\mu(\tau_{t},t)-\frac{1}{2}\sigma(t)^{2}\nabla_{\tau_{t}}\log p_{t}(\tau_{t})\right]dt$. In score-based diffusion models, the score function, $\nabla_{\tau_{t}}\log p_{t}(\tau_{t},t)$, is typically parametrized by a neural network $s_{\theta}(\tau,t)$. The model is trained with a score-matching loss: $\min_{\theta}\mathbb{E}_{p(\tau_{t},t)}\left[\left\|s_{\theta}(\tau_{t},t)-\nabla_{\tau_{t}}\log p(\tau_{t},t)\right\|_{2}^{2}\right]$.
The formulation provided for score-matching diffusion above is of the continuous Stochastic Differential Equation (SDE) form, where the denoising timesteps are represented in continuous time. However, the SDE can be discretized to retrieve a discrete form that matches the Denoising Diffusion Probabilistic Models (DDPM) [35] formulation commonly referred to as the Variance Preserving (VP) SDE [36]. In this formulation, the forward noising process is defined as: $\tau_{i}=\sqrt{\alpha_{i}}\tau_{i-1}+\sqrt{1-\alpha_{i}}z,z\sim\mathcal{N}(0,I)$. Following the convention from DDPM, $\beta_{i}$ represent the diagonal matrix of variances parameterizing the normal distribution, $\alpha_{i}=1-\beta_{i}$, and $\bar{\alpha}_{i}=\prod_{s=0}^{i}\alpha_{s}$. The variance parameter $\beta_{i}$ is defined by the noise variance schedule, where we use the standard cosine schedule. Complete noise is defined at $i=T$ while the data distribution is defined at $i=0$. The relationship between the score-function $s_{\theta}(\tau_{i},i)$ and the model learned through DDPM’s epsilon matching $z_{\theta}(\tau_{i},i)$ is $z_{\theta}(\tau_{i},i)=-\sqrt{1-\bar{\alpha}_{i}}s_{\theta}(\tau_{i},i)$. For brevity, we omit $i$ and write $s_{\theta}(\tau_{i})$ rather than $s_{\theta}(\tau_{i},i)$.
@@ V-A Kinematic Constraint Gradient Guidance
In this subsection, we outline the standard formulation for sampling from diffusion models without incorporating motion planning constraints. Then, we describe prior work for sampling with constraints and introduce our novel sampling method. We refer to our method as Kinematic Constraint Gradient Guidance (KCGG).
@@
First, the standard sampling procedure for the diffusion model without constraints conducts one denoising step with Equation 1.
$\tau_{i-1}=\frac{1}{\sqrt{\alpha_{i}}}\left(\tau_{i}+\left({1-\alpha_{i}}\right)s_{\theta}(\tau_{i})\right)+\sqrt{\sigma_{i}}z,z\sim\mathcal{N}(0,1)$ (1)
Here, $s_{\theta}$ denotes the score function parameterized by the diffusion model and $z$ is Gaussian noise parameterized by the denoising process.
@@
In order to generate trajectory samples conditionally with respect to the optimality condition $p(\tau|\mathcal{O})$, prior works like Motion Planning Diffusion (MPD1) [24] and Diffuser [23] choose to directly guide the noisy trajectories $\tau_{i-1}$.
$\tau_{i-1}^{\prime}$ $=\frac{1}{\sqrt{\alpha_{i}}}\left(\tau_{i}+\left({1-\alpha_{i}}\right)s_{\theta}(\tau_{i})\right)$ (2)
$\tau_{i-1}$ $=\tau_{i-1}^{\prime}-\mathbf{\nabla_{\tau_{i-1}^{\prime}}c_{j}(\tau_{i-1}^{\prime})}+z,z\sim\mathcal{N}(0,I)$
One problem with this formulation is that the constraint projects the intermediate diffusion samples out of distribution of the diffusion model’s trained denoising paths [47]. We refer the reader to [48] for proofs on the coherence of these projection-based methods for diffusion model sampling.
@@
KCGG modifies the diffusion sampling procedure to Equation 3.
$\tau_{i-1}$ $=\tau_{i-1}^{\prime}-\mathbf{\nabla_{\tau_{i}}c_{j}(\hat{\tau}_{0})}+z,z\sim\mathcal{N}(0,I)$ (3)
@@
By computing the gradient of the constraint for the predicted clean trajectory $c_{j}(\hat{\tau_{0}})$ rather than a noisy trajectory $c_{j}(\tau_{i-1}^{\prime})$, the constraints can be better informed. Crucially, the cost functions $c_{j}$ are not defined for noisy trajectories but rather noiseless trajectories. Therefore by using $\hat{\tau}_{0}$ we get a better gradient from the cost function. We can approximate $\hat{\tau_{0}}$ via $\tau_{i-1}$, as shown in Equation 4.
@@
Second, we compute the gradient with respect to $\tau_{i}$ rather than $\tau_{i-1}$. The trajectory $\tau_{i}$ is the noisy trajectory before the denoising step. The reason we choose $\tau_{i}$ rather than $\tau_{i-1}$ is that we derive our procedure as sampling from the posterior distribution of $p(\tau|c)$. As we define our process through a score-based approach, we are looking to compute $\nabla_{\tau_{i}}\log p(\tau|c)$. Using Bayes rule, we therefore can compute the posterior with $\nabla_{\tau_{i}}\log p(\tau_{i}|c)=\nabla_{\tau_{i}}\log p(c|\tau_{i})+\nabla_{\tau_{i}}\log p(\tau_{i})$.
Given that $\hat{\tau_{0}}$ is a function of $\tau_{i}$, our approach takes the gradient not only through the constraint function, $c_{j}$ as prior methods did but also through the diffusion model, $s_{\theta}$ itself. Intuitively, by taking the gradient through both the constraint function $c_{j}$ and the diffusion model $s_{\theta}$, we modify the entire trajectory $\tau_{i-1}$ to adhere to the constraint rather than just a portion of the trajectory. This whole-trajectory update is generally superior to updating small portions as it ensures the entire trajectory remains coherent and smooth, with all timesteps adjusted in relation to each other. It allows for global optimization of the trajectory, potentially finding better solutions that satisfy the constraint while maintaining overall motion quality. Furthermore, by updating the whole trajectory, we reduce the risk of getting stuck in local minima that may occur when only modifying small portions.
Many motion planning task constraints, such as constraining the start or ending states, only act on a few timesteps within the trajectory. Similarly, our constraints only enforce the end-effector to be at the ball at a single timestep. In traditional approaches, the function $\nabla_{\tau_{i-1}}c(\tau_{i-1})$ only updates a small portion of the predicted trajectory at each denoising timestep, where all other points receive no updates as the gradient is 0. This limited update can lead to suboptimal solutions, as it doesn’t consider the impact on the entire motion sequence. In contrast, our method’s whole-trajectory update allows for more flexible and globally optimal solutions, even when constraints are localized to specific timesteps.

### v2

@@ Abstract
Advances in robot learning have enabled robots to generate skills for a variety of tasks. Yet, robot learning is typically sample inefficient, struggles to learn from data sources exhibiting varied behaviors, and does not naturally incorporate constraints. These properties are critical for fast, agile tasks such as playing table tennis. Modern techniques for learning from demonstration improve sample efficiency and scale to diverse data, but are rarely evaluated on agile tasks. In the case of reinforcement learning, achieving good performance requires training on high-fidelity simulators. To overcome these limitations, we develop a novel diffusion modeling approach that is offline, constraint-guided, and expressive of diverse agile behaviors. The key to our approach is a kinematic constraint gradient guidance (KCGG) technique that computes gradients through both the forward kinematics of the robot arm and the diffusion model to direct the sampling process. KCGG minimizes the cost of violating constraints while simultaneously keeping the sampled trajectory in-distribution of the training data. We demonstrate the effectiveness of our approach for time-critical robotic tasks by evaluating KCGG in two challenging domains: simulated air hockey and real table tennis. In simulated air hockey, we achieved a 25.4% increase in block rate, while in table tennis, we achieved a 17.3% increase in success rate compared to imitation learning baselines.
@@ I Introduction
Figure 1: Overview of KCGG: At inference, KCGG computes the cost given the constraint and intermediate trajectory at each denoising step. The gradient with respect to the prior trajectory through the diffusion model is used to update the sample.
In this paper, we develop a novel imitation learning based approach that enables robots to learn from small offline datasets to perform agile tasks in dynamical environments. We ground our work in an application to striking motions evaluated in established challenging domains for robotics: a virtual air hockey domain [20, 9] and a physical table tennis domain [21, 22]. We show that our diffusion-based approach can learn to condition trajectories to meet environment constraints and perform tasks difficult to specify (e.g. hitting a topspin, a high lob, etc) with just 15 kinesthetic demonstrations per stroke type, which classical motion planning cannot achieve without explicit task heuristics. The key to our approach is the novel formulation of a kinematic constraint gradient guidance (KCGG) technique that enables the robot to balance between constraints and adhering to the demonstration distribution.
@@
  • We show KCGG improves a real-world table tennis robot’s ability to return ping-pong balls. Our method improves success rate by 124.5% over diffusion baselines [23, 24] and 17.3% over state-of-the-art imitation learning baselines [19], demonstrating efficacy in agile robotic tasks.
@@ II Related Work
LfD [15, 16, 17] aims to enable robots to acquire new skills by learning to imitate an expert. MP-based LfD has been explored extensively for agile striking motions in table tennis by learning movement primitives. For example, Mixture of Movement Primitives (MoMP) [22, 25] learns to select a dynamic movement primitive (DMP) from a library of pre-trained DMPs to generate joint trajectories for ball returns. Probabilistic movement primitives (ProMP) [26], which model the variability in movements as trajectory distributions, were adapted to generate striking motions that maximize the likelihood of ball return [27, 28, 29, 30]. Movement primitive approaches are sample-efficient; however, generating diverse motions (e.g. slice vs. lob in table tennis) is difficult without a hierarchical approach.
Recently, RL methods are enabling systems to explore in task space without using demonstrated expert knowledge [21, 31, 32, 33, 34]. However, most of them require large training datasets [21, 32], carefully shaped rewards [21], or human behavior modeling [34] which require a simulation. In contrast, we develop a fully offline method that requires few demonstrations and captures diverse agile behaviors in a single model, simplifying the training and deployment process onto real-world robots.
@@
Diffusion models [35, 36] are a class of generative models that have recently shown promise in robotics for imitation learning [23], offline RL [37, 38], and motion planning [23, 39]. Compared with previous generative modeling approaches such as GANs [40], diffusion models generally exhibit improved sample quality and diversity with a tradeoff in sampling time. In this work, we enhance the quality of samples generated by diffusion models while maintaining similar sampling time. This approach produces higher-quality samples without increasing computational cost, which is crucial for learning agile motions.
Recent works using diffusion models for robotic manipulation, such as BESO [41] and MPD2 [42], demonstrated impressive performance and high success rates in challenging domains requiring millimeter-level precision. Our work, however, explores the additional complexity of solving dynamic tasks. Furthermore, while BESO [41] is designed for goal-conditioned tasks, the goal space in our domain is difficult to define. Consequently, our method does not rely on goal conditioning and instead adapts during sampling time.
One advantage of diffusion models is the ability to guide samples towards certain objectives at sampling time. Several prior works in robotics have shown how to add constraints for motion planning [24] and state estimation [43]. The computer vision community has also investigated using diffusion models to solve inverse problems [44, 45], where the goal is to reconstruct samples from sparse measurements. We take inspiration from these techniques used in inverse problems to formulate a novel method for guiding diffusion samples towards kinematics constraints and evaluate this method in a difficult agile robotics domain. Our method is simple and enables the constraints to be better incorporated into the resulting diffusion samples.
Figure 2: Visualization of Diffusion Sampling Process in Air-Hockey Defend: KCGG gradually shifts the conditioned points in the trajectory (shown in green) towards a predicted future puck position (shown as the orange point).
@@ III Problem Formulation
The agile robot striking problem can be formulated as a motion planning task, where the objective is to determine a trajectory for the robot that adheres to a variety of constraints. We define $\tau=\{s_{0},s_{1},...s_{H}\}$ as a sequence of states up to a time horizon, $H$. Each state $s=[q,\dot{q}]\in\mathbb{R}^{d}$ is composed of the robot’s joint positions, $q$, and joint velocities, $\dot{q}$. The goal of motion planning is to identify a trajectory between initial and final states that satisfy a set of costs or constraints, $c_{j}(\tau)$. The objective is to minimize the total cost $\mathcal{J}(\tau)=\sum_{j}\lambda_{j}c_{j}(\tau)$, subject to kinematic or other constraints. Other approaches formulate the motion-planning problem through probabilistic inference. Here, the optimality of a state within the trajectory is represented as $\mathcal{O}_{t}$ and the goal is to sample from a posterior distribution, $p(\tau|\mathcal{O}_{1:t})\sim p(\tau)p(\mathcal{O}_{1:t}|\tau)$ [23]. We use this formulation, where the diffusion model learns the prior, $p(\tau)$, and we use the guided sampling formulation [46] to conditionally sample the model on our task constraints.
A task constraint of interest for the agile striking problem is to ensure contact between the end-effector and the moving object, where $b\in\mathbb{R}^{3}$ is the object’s position in Cartesian space. The forward kinematics function that maps the robot joint positions, $q$, to end-effector position, $x$, is denoted as $F\colon q\rightarrow x,q\in R^{d},x\in R^{3}$. Our constraint minimizes the distance between the end-effector and object positions at a timestep, $t$, within a time window, $[t_{s},t_{e}]$, where the robot can feasibly strike the object, i.e. $c(\tau)=\min_{t\in[t_{s},t_{e}],q_{t}\in\tau}\|F(q_{t})-b_{t}\|_{2}^{2}$. Our method is amenable to other types of constraints (e.g. for obstacle avoidance).
@@ IV Preliminaries: Score-Based Diffusion Models
Diffusion models are a class of generative models that learn to generate samples through a forward noising process and a reverse de-noising process. The forward noising process is commonly defined by: $d\tau_{t}=f(\tau_{t},t)\,dt+\sigma(t)\,dw_{t}$, where $\tau$ is the state trajectory of the robot, $f(\tau_{t},t)$ is the drift term and $\sigma(t)$ is the standard Brownian noise process. The forwards process is coupled with a backwards process: $d\tau_{t}=\left[\mu(\tau_{t},t)-\frac{1}{2}\sigma(t)^{2}\nabla_{\tau_{t}}\log p_{t}(\tau_{t})\right]dt$. In score-based diffusion models, the score function, $\nabla_{\tau_{t}}\log p_{t}(\tau_{t},t)$, is typically parametrized by a neural network, $s_{\theta}(\tau,t)$. The model is trained with a score-matching loss: $\min_{\theta}\mathbb{E}_{p(\tau_{t},t)}\left[\left\|s_{\theta}(\tau_{t},t)-\nabla_{\tau_{t}}\log p(\tau_{t},t)\right\|_{2}^{2}\right]$.
The formulation provided for score-matching diffusion above is of the continuous Stochastic Differential Equation (SDE) form, where the denoising timesteps are represented in continuous time. However, the SDE can be discretized to retrieve a discrete form that matches the Denoising Diffusion Probabilistic Models (DDPM) [35] formulation commonly referred to as the Variance Preserving (VP) SDE [36]. In this formulation, the forward noising process is defined as: $\tau_{i}=\sqrt{\alpha_{i}}\tau_{i-1}+\sqrt{1-\alpha_{i}}z,z\sim\mathcal{N}(0,I)$. Following the convention from DDPM, $\beta_{i}$ represent the diagonal matrix of variances parameterizing the normal distribution, $\alpha_{i}=1-\beta_{i}$, and $\bar{\alpha}_{i}=\prod_{s=0}^{i}\alpha_{s}$. The variance parameter, $\beta_{i}$, is defined by the noise variance schedule, where we use the standard cosine schedule. Complete noise is defined at $i=T$ while the data distribution is defined at $i=0$. The relationship between the score-function, $s_{\theta}(\tau_{i},i)$, and the model learned through DDPM’s epsilon matching, $z_{\theta}(\tau_{i},i)$, is $z_{\theta}(\tau_{i},i)=-\sqrt{1-\bar{\alpha}_{i}}s_{\theta}(\tau_{i},i)$. For brevity, we omit $i$ and denote $s_{\theta}(\tau_{i})$ as $s_{\theta}(\tau_{i},i)$.
@@ V-A Kinematic Constraint Gradient Guidance
In this subsection, we outline the standard formulation for sampling from diffusion models without incorporating motion planning constraints. Then, we describe prior work for sampling with constraints and introduce our novel sampling approach.
@@
Each denoising step in a standard, unconstrained diffusion model follows Eq. 1, where $s_{\theta}$ denotes the score function parameterized by the diffusion model and $z\sim\mathcal{N}(0,1)$ is Gaussian noise parameterized by the denoising process.
$\tau_{i-1}=\frac{1}{\sqrt{\alpha_{i}}}\left(\tau_{i}+\left({1-\alpha_{i}}\right)s_{\theta}(\tau_{i})\right)+\sqrt{\sigma_{i}}z$ (1)
@@
To generate trajectory samples conditionally with respect to the optimality condition, $p(\tau|\mathcal{O})$, prior works choose to directly guide the noisy trajectories, $\tau_{i-1}$, as follows in Eq. 2.
$\tau_{i-1}^{\prime}$ $=\frac{1}{\sqrt{\alpha_{i}}}\left(\tau_{i}+\left({1-\alpha_{i}}\right)s_{\theta}(\tau_{i})\right),$ (2)
$\tau_{i-1}$ $=\tau_{i-1}^{\prime}-\mathbf{\nabla_{\tau_{i-1}^{\prime}}c_{j}(\tau_{i-1}^{\prime})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{\sigma_{i}}}z$
However, with this formulation, the constraint projects the intermediate diffusion samples out of distribution from the diffusion model’s trained denoising paths [47]. We refer the reader to [48] for proofs on the coherence of these projection-based methods for diffusion model sampling.
@@
KCGG modifies the diffusion sampling procedure to Eq. 3.
$\tau_{i-1}$ $=\tau_{i-1}^{\prime}-\mathbf{\nabla_{\tau_{i}}c_{j}(\hat{\tau}_{0})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{\sigma_{i}}}z$ (3)
@@
By computing the gradient of the constraint for the predicted clean trajectory, $c_{j}(\hat{\tau_{0}})$, rather than a noisy trajectory, $c_{j}(\tau_{i-1}^{\prime})$, the constraints can be better informed. Crucially, the cost functions, $c_{j}$, are not defined for noisy trajectories but rather noiseless trajectories. Therefore by using $\hat{\tau}_{0}$ we get a better gradient from the cost function. We can approximate $\hat{\tau_{0}}$ via $\tau_{i-1}$, as shown in Equation 4.
@@
Second, we compute the gradient with respect to $\tau_{i}$ rather than $\tau_{i-1}$. $\tau_{i}$ is the noisy trajectory before the denoising step. We choose $\tau_{i}$ rather than $\tau_{i-1}$ because we derive our procedure as sampling from the posterior distribution of $p(\tau|c)$. As we define our process through a score-based approach, we compute $\nabla_{\tau_{i}}\log p(\tau|c)$. Using Bayes rule, we therefore can compute the posterior with $\nabla_{\tau_{i}}\log p(\tau_{i}|c)=\nabla_{\tau_{i}}\log p(c|\tau_{i})+\nabla_{\tau_{i}}\log p(\tau_{i})$.
@@
Given that $\hat{\tau_{0}}$ is a function of $\tau_{i}$, our approach takes the gradient not only through the constraint function, $c_{j}$, as prior methods did, but also through the diffusion model, $s_{\theta}$. Intuitively, by taking the gradient through both $c_{j}$ and $s_{\theta}$, the entire trajectory, $\tau_{i-1}$, adheres to the constraint rather than just a portion of the trajectory. This whole-trajectory update is generally superior to updating small portions as it ensures the entire trajectory remains coherent and smooth. Furthermore, by updating the whole trajectory, we reduce the risk of getting stuck in local minima that may occur when only modifying small portions.
Many motion planning task constraints, e.g. constraints on goal states, only act on a subset of the trajectory. In prior approaches, the function, $\nabla_{\tau_{i-1}}c(\tau_{i-1})$ only updates the constrained subset of the predicted trajectory at each denoising step. However, this can lead to suboptimal solutions, as it does not consider the impact on the entire motion sequence. In contrast, our method’s whole-trajectory update allows for more flexible and globally optimal solutions, even when constraints are localized to specific timesteps.